### v2

# Memory-Efficient Personalization of Text-to-Image Diffusion Models via Selective Optimization Strategies

Abstract
Memory-efficient personalization is critical for adapting text-to-image diffusion models while preserving user privacy and operating within the limited computational resources of edge devices. To this end, we propose a selective optimization framework that adaptively chooses between backpropagation on low-resolution images (BP-low) and zeroth-order optimization on high-resolution images (ZO-high), guided by the characteristics of the diffusion process. As observed in our experiments, BP-low efficiently adapts the model to target-specific features, but suffers from structural distortions due to resolution mismatch. Conversely, ZO-high refines high-resolution details with minimal memory overhead but faces slow convergence when applied without prior adaptation. By complementing both methods, our framework leverages BP-low for effective personalization while using ZO-high to maintain structural consistency, achieving memory-efficient and high-quality fine-tuning. To maximize the efficacy of both BP-low and ZO-high, we introduce a timestep-aware probabilistic function that dynamically selects the appropriate optimization strategy based on diffusion timesteps. This function mitigates the overfitting from BP-low at high timesteps, where structural information is critical, while ensuring ZO-high is applied more effectively as training progresses. Experimental results demonstrate that our method achieves competitive performance while significantly reducing memory consumption, enabling scalable, high-quality on-device personalization without increasing inference latency.
[…]
1 Introduction
Figure 1: Key observations on different optimization strategies. (a) ZO-high requires target-specific knowledge; applying it from scratch fails to personalize the model, while using it after partial fine-tuning improves personalization. (b) Comparison of fine-tuning results across resolutions and timestep ranges using backpropagation. BP-low suffers from structural distortions due to resolution mismatch between training and inference, especially in higher timestep ranges, while BP-high achieves strong personalization but is memory-intensive and impractical for edge devices.
The increasing popularity of text-to-image diffusion models has fueled demand for personalized content generation, enabling applications such as avatar creation [17, 7] and object customization in novel contexts [2]. To meet this demand, training-free personalization methods [16, 9, 15, 18] have emerged, eliminating the need for test-time fine-tuning. These approaches streamline personalization during inference; however, they often incur high memory usage, increased latency, or substantial pre-training requirements, which limit their practicality for on-device deployment.
In contrast, fine-tuning-based methods such as DreamBooth [14], Textual Inversion [4], and Custom Diffusion [8] adapt model weights using user-provided images, enabling high-fidelity personalization with fine-grained details. Due to their high computational demands, these methods are typically executed on servers. However, growing privacy concerns have increased interest in on-device personalization, which allows users to fine-tune models locally without transmitting sensitive data. While this approach enhances privacy, the fine-tuning process remains computationally intensive and memory-demanding, posing challenges for deployment on resource-constrained edge devices.
To address this, parameter-efficient fine-tuning techniques such as LoRA [6] have been explored, which reduce the number of trainable parameters through low-rank adaptation. Although effective, LoRA still relies on full backpropagation, which requires storing activations and gradients, resulting in significant memory overhead. Zeroth-order optimization methods such as MeZO [10] avoid backpropagation by estimating gradients through forward passes, offering a more memory-efficient alternative.
However, applying MeZO directly to a pretrained model leads to poor gradient estimates and slow convergence due to the absence of target-specific cues. We observed that MeZO becomes significantly more effective when preceded by partial fine-tuning that introduces target-specific information (Observation 1), as shown in Fig. 1 (a). To provide such information, high-resolution backpropagation would be ideal, but it is too memory-intensive for edge devices. A practical alternative is to use low-resolution images (BP-low), which reduces memory usage by lowering activation size. Nevertheless, BP-low struggles to capture the global structure of the target object and often produces distorted outputs (Observation 2), particularly in the higher timestep range (750–1000), where structural information is more critical, as illustrated in Fig. 1 (b).
To overcome these limitations, we propose a selective optimization framework that combines BP-low for lightweight adaptation with zeroth-order optimization on high-resolution images (ZO-high) for structural refinement. While BP-low efficiently injects target-specific features, it tends to distort global structure at high timesteps. In contrast, ZO-high directly operates on high-resolution images, which helps preserve structural details but remains less effective early in training due to weak gradient estimates. To address this, we introduce a timestep-aware probabilistic function that dynamically selects between the two methods based on the diffusion timestep and training progress, supporting efficient personalization on edge devices without altering the standard sampling process.
Our contributions are summarized as follows:
  • We propose a novel framework that combines low-resolution backpropagation and high-resolution zeroth-order optimization for memory-efficient personalization.
  • We introduce a timestep-aware probabilistic function that dynamically selects the optimization strategy based on diffusion timestep and training progress.
  • We demonstrate that our method enables scalable, high-quality on-device personalization with significantly reduced memory usage.
2 Proposed Method
We propose a memory-efficient optimization framework for on-device personalization of text-to-image diffusion models. Our method addresses the limitations of low-resolution backpropagation and Memory-efficient Zeroth-order Optimizer (MeZO) [10] by dynamically selecting either technique based on the diffusion timestep and training process. This hybrid strategy enables high-quality fine-tuning with low memory usage, making it suitable for edge deployment.
The proposed framework uses backpropagation on low-resolution images (BP-low) to inject target-specific information efficiently, and zeroth-order optimization on high-resolution images (ZO-high) to refine global structure with low memory overhead. A timestep-aware probabilistic function selects the method at each step, reducing BP-low overfitting and structural distortions. The full training procedure is summarized in Algorithm 1.
2.1 Overview of the Optimization Framework
Given personalization images $\mathcal{X}$ and a prompt $c$ (e.g., “a scs dog”), the goal is to fine-tune a pretrained diffusion model $\theta$ to generate images that reflect user-specific content.
At each training step $i$, an image $x$ is sampled from $\mathcal{X}$, and a diffusion timestep $t$ is drawn from a uniform distribution $t\sim\mathcal{U}(0,T)$. Based on $t$ and $i$, a probabilistic function selects either BP-low or ZO-high.
2.2 Backpropagation on Low-Resolution Images
Backpropagation is applied to low-resolution images $x_{\text{low}}$ (e.g., $256\times 256$) to reduce memory usage while capturing coarse personalization features. The optimization follows the standard gradient descent update:
$\theta_{i+1}=\theta_{i}-\eta\nabla_{\theta}\mathcal{L}_{\text{bp}}(\theta_{i};x_{\text{low}},c,t),$ (1)
where $\theta_{i}$ denotes the parameters at training step $i$, $\eta$ is the learning rate, and $\mathcal{L}_{\text{bp}}$ is the diffusion reconstruction loss at timestep $t$.
2.3 MeZO on High-Resolution Images
We use MeZO with gradient accumulation to improve gradient approximation while maintaining memory efficiency. For each original image $x$ (e.g., $512\times 512$), MeZO computes the accumulated gradient. For each perturbation $n\in\{1,2,\dots,N\}$, we sample a random perturbation vector $\boldsymbol{z}^{(n)}\sim\mathcal{N}(0,I_{d})$ from a standard normal distribution and compute the gradient estimate:
$\hat{g}_{i}^{(n)}=\frac{\mathcal{L}(\theta_{i}+\epsilon\boldsymbol{z}^{(n)};x,c,t)-\mathcal{L}(\theta_{i}-\epsilon\boldsymbol{z}^{(n)};x,c,t)}{2\epsilon}\boldsymbol{z}^{(n)}.$ (2)
The final gradient estimate $\hat{g}_{i}$ is obtained by a weighted sum of the $N$ estimates, $\hat{g}_{i}=\sum_{n=1}^{N}w_{n}\hat{g}_{i}^{(n)}$, where $w_{n}=\frac{1}{N}$. The parameter update is given by $\theta_{i+1}=\theta_{i}-\alpha\hat{g}_{i}$, with $\alpha$ as the learning rate for MeZO updates.
Figure 2: Comparison of TAP and DTAP over training steps when $t_{\text{max}}=1000$, $i_{\text{max}}=1000$, $k=0.05$, and $t_{\text{mid}}=750$. The black dashed line represents the Timestep-Aware Probability (TAP), where $t_{\text{mid}}$ remains fixed throughout training. In contrast, the colored curves illustrate the Dynamic Timestep-Aware Probability (DTAP) at different training steps ii.
2.4 Timestep-Aware Probabilistic Function
A key component of our framework is the probabilistic function that selects BP-low or ZO-high at each training step. This function adapts the optimization strategy based on the diffusion timestep $t$ and training step $i$, helping to reduce overfitting and structural artifacts.
2.4.1 Timestep-Aware Probability (TAP)
Following Observation 2, BP-low tends to distort global structure at high timesteps ($t\approx t_{\text{max}}$), where structural information is critical [3]. To address this, we define the probability of selecting ZO-high as:
$p^{zo}_{t}=\frac{1}{1+e^{-k(t-t_{\text{mid}})}},$ (3)
where $k>0$ controls the steepness, and $t_{\text{mid}}$ is the midpoint of the transition. This ensures ZO-high is more likely to be selected at higher timesteps (see Fig. 2).
2.4.2 Dynamic Timestep-Aware Probability (DTAP)
To incorporate Observation 1, which states that MeZO becomes more effective as training progresses, we extend TAP to dynamically adjust the midpoint over time. Specifically, we define:
$t_{\text{dyn}}(i)=t_{\text{start}}+\frac{i}{i_{\text{max}}}(t_{\text{end}}-t_{\text{start}}),$ (4)
where $t_{\text{start}}=t_{\text{max}}$ and $t_{\text{end}}=2t_{\text{mid}}-t_{\text{max}}$. At $i=0.5i_{\text{max}}$, $t_{\text{dyn}}(i)=t_{\text{mid}}$, aligning DTAP with TAP. The final probability becomes:
$p^{zo}_{i,t}=\frac{1}{1+e^{-k(t-t_{\text{dyn}}(i))}}.$ (5)
This adaptive selection enables the framework to leverage the strengths of both BP-low and ZO-high throughout training, improving personalization quality while maintaining memory efficiency.
Algorithm 1 Training Algorithm
1:Pretrained diffusion model $\theta_{0}$, personalization images $\mathcal{X}$, prompt $c$, total number of training steps $i_{\text{max}}$, maximum timestep $t_{\text{max}}$, number of perturbations NN
2:for training step $i=1$ to $i_{\text{max}}$ do
3:  Sample image $x\in\mathcal{X}$
4:  Sample diffusion timestep $t\sim\mathcal{U}(0,t_{\text{max}})$
5:  Compute $p^{zo}_{i,t}$ using Eq. 5
6:  if random() $p^{zo}_{i,t}$ then $\triangleright$ BP-low
7:   Downsample $x$ to $x_{\text{low}}$
8:   Compute gradient $\nabla_{\theta}\mathcal{L}_{\text{bp}}(\theta_{i};x_{\text{low}},c,t)$
9:   Update $\theta_{i+1}=\theta_{i}-\eta\nabla_{\theta}\mathcal{L}_{\text{bp}}$
[…]
11:   Use high-resolution image $x$
12:   Initialize accumulated gradient $\hat{g}_{i}=0$
13:   for $n=1$ to NN do
14:     Sample perturbation $\boldsymbol{z}^{(n)}$
15:     Compute $\hat{g}_{i}^{(n)}$ using Eq. 2
16:     Accumulate $\hat{g}_{i}=\hat{g}_{i}+w_{n}\hat{g}_{i}^{(n)}$
17:   end for
18:   Update $\theta_{i+1}=\theta_{i}-\alpha\hat{g}_{i}$
19:  end if
20:end for
Figure 3: Qualitative comparison of optimization methods. Overfitting (BP-high), distortions (BP-low), poor personalization (ZO-high), instability (Random), and artifacts (Reversed scheduling) are observed. Our method maintains subject fidelity and structural consistency.
3 Experiments
3.1 Experimental Setup
We evaluate our method on multiple text-to-image diffusion models, including SD V1.5, SD V2.1 [13], SDXL [11], and SSD-1B [5], focusing on models suitable for edge-device deployment. All models were fine-tuned based on DreamBooth-LoRA [14], with a fixed rank of 4. We adopt DTAP as our optimization selection strategy and use the hyperparameters specified in Fig. 2. Experiments were conducted on the DreamBooth dataset [14], which includes 30 subjects and 25 prompts per subject. We report CLIP-I, CLIP-T [12], and DINO [1] scores to assess subject similarity and text-image alignment.
Method Performance Metrics Mem (MiB)
DINO CLIP-I CLIP-T BP ZO
BP-High 0.6434 0.7926 0.2890 5722 -
Ours (r=0.750) 0.6403 0.7921 0.2904 5033 4756
Ours (r=0.625) 0.6433 0.7884 0.2914 4737 4756
SD V1.5 Ours (r=0.500) 0.6437 0.7885 0.2904 4547 4756
BP-High 0.6762 0.8075 0.2961 6485 -
Ours (r=0.750) 0.6758 0.8058 0.2969 5832 5613
Ours (r=0.625) 0.6787 0.8043 0.2955 5561 5613
SD V2.1 Ours (r=0.500) 0.6833 0.8067 0.2938 5368 5613
BP-High 0.7329 0.8121 0.3002 14397 -
Ours (r=0.750) 0.7269 0.8084 0.3003 11260 9546
Ours (r=0.625) 0.7373 0.8141 0.2957 9958 9546
SDXL Ours (r=0.500) 0.7336 0.8135 0.2941 8957 9546
BP-High 0.6899 0.7955 0.3069 9030 -
Ours (r=0.750) 0.6913 0.7985 0.3048 7115 7109
Ours (r=0.625) 0.7112 0.8048 0.3022 6338 7109
SSD-1B Ours (r=0.500) 0.7080 0.8075 0.2996 5729 7109
Table 1: Quantitative comparison across different diffusion models. Our method achieves comparable or superior performance to BP-high while reducing memory usage. Peak memory is defined as the maximum of BP and ZO during training and is underlined.
3.2 Quantitative Performance
We assess the scalability of our method across various diffusion models and compare it with full-resolution backpropagation. To analyze the effect of resolution, we vary the resize ratio $r\in\{0.5,0.625,0.75\}$. As shown in Tab. 1, our method consistently matches or outperforms BP-high while significantly reducing memory usage. Notably, on SDXL, it achieves improvements across all metrics with up to 33.69% lower memory consumption. Here, peak memory is defined as the maximum of BP and ZO memory usage during training. These results demonstrate the effectiveness of combining low-resolution backpropagation and zeroth-order optimization with a timestep-aware selection strategy.
3.3 Qualitative Performance
Fig. 3 presents qualitative comparisons across different optimization strategies, conducted using SD 2.1. BP-low introduces structural distortions, particularly at high timesteps. ZO-high fails to capture target-specific details due to slow convergence. Random selection between BP-low and ZO-high results in unstable training and inconsistent personalization. Setting the steepness parameter $k$ to a negative value results in a reversed scheduling, where ZO-high is applied in the early stages of training and BP-low is deferred to higher timesteps. This configuration leads to ineffective personalization and increased structural distortions, highlighting the importance of our intended timestep-aware scheduling. In contrast, our method maintains structural consistency while preserving subject-specific features. In some cases, it even surpasses BP-high in text alignment, suggesting improved generalization and reduced overfitting.
For further details, including implementation information, extended comparisons, ablation studies, hyperparameter selection, user study results, and related works, please refer to the supplementary material.
4 Conclusion
We presented a memory-efficient optimization framework for on-device personalization of text-to-image diffusion models. Rather than naively combining low-resolution backpropagation and zeroth-order optimization, our method adaptively selects between them based on key observations of diffusion timesteps. This improves personalization and structural consistency, and highlights the potential of resolution-aware training and gradient-free optimization for diffusion models. We hope this encourages further exploration of efficient personalization strategies.